OpenStudy (anonymous):

The sum of five times a number and four is equal to negative eleven. What is the number? -2 -3 -4 -5

OpenStudy (anonymous):

maybe -3?

OpenStudy (anonymous):

for x?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

5x+4=-11 -4=-4 5x=-15 5x/5=-15/5 x=-3 check plug in the number do order of operation 5*-3+4=-11 -15+4=11

OpenStudy (anonymous):

so you tell me now! :D

OpenStudy (anonymous):

yes, -3

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

so it is -3 then

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i spaced out -3 on purpose